# Assessing the viability of non-light water reactor concepts for electricity and heat generation in decarbonized energy systems

### v1

@@ 5 Discussion
The results shown above underline the lack of cost competitiveness of nuclear power compared to other energy generation technologies. They show that the cost-efficient capacity expansion of especially non-light water reactors would require substantial cost reductions that are, as described in Section 3.3, not to be expected in the coming years.
@@
However, if the above-mentioned optimistic cost reductions were to be achieved, the results show that non-light water reactor technologies could contribute to energy system decarbonization in Europe in this particular setting (e.g., assuming favorable capacity factors and industry capability, etc.). However, this is limited to mostly HTRs, which provide valuable heat energy at temperatures of up to 1000°C.
@@
But regardless of costs, nuclear reactors feature additional challenges that would need to be overcome for a substantial capacity ramp-up. For example, the heterogeneity of fuel requirements for different SNR concepts corresponds with that of the reactor designs themselves. Many concepts additionally increase proliferation risks because of their requirement of high assay low enriched uranium (HALEU) fuel for which specialized fuel production facilities would have to be funded and constructed-which in turn requires substantial investment and long-term planning and the limitation of industry efforts to a handful of promising reactor designs [120, 121]. It is uncertain whether this type of infrastructure could be up and running in time for these reactor concepts to play a significant role in decarbonizing the European energy system.
@@
Furthermore, the cost data mentioned above do not include external costs of nuclear, such as decommissioning and waste management, or social costs, such as risks relating to accidents, proliferation, or uranium mining and tilling [122, 123, 124, 125]. Additional waste challenges arise from implementing non-light water reactors that generate different kinds of waste from, e.g., HALEU fuel, creating additional and currently unstudied waste streams that would have to be dealt with [121]. However, with most nuclear waste responsibility lying with the respective governments [126, 127], substantial funding uncertainties could be removed from projects [128], as suggested to be implemented by Sawicki and Horbaczewska [129]. This would, however, require substantial political power to be achieved, especially in Europe. Additionally, long-term financial assurance and commitment to a limited number of reactor designs are necessary [17].
@@
Additionally, the assumed flexibility of nuclear power plant operations without constraints can be contested. Lynch et al. [116] demonstrate that today’s LWRs can be an effective technology for load following, and Jenkins et al. [19] show in their model for the South-West U.S. that flexible nuclear operations can reduce system costs and increase reactor operator revenues. Loisel et al. [130, p. 281] even claim that "in practice, countries with […] high intermittent renewables […] need [nuclear power plants] to operate". However, despite theoretical technical feasibility, current nuclear reactors’ ability to run flexibly is seldom applied today, with operational experience limited mostly to France [42]. In the U.S., operational experience of flexibly dispatchable nuclear power plants has been limited despite regulations allowing for flexible operations and load following [131]. Reactors designed to operate as flexible backup capacities in energy systems based mainly on renewables face challenges of long standstills and fast ramp-ups. Running at high capacity factors generating heat for other purposes and then flexibly switching to power generation when required will be technologically challenging to implement, least finance [132, 51].
@@
Our model approach faces some inevitable limitations that could limit the applicability of the results. First, the potential locality of heat demand is not accounted for. This is relevant because, for direct heat provision, sources must be located close to off-takers to minimize losses during transfer. This could be especially relevant for nuclear reactors as these technologies may face substantial opposition from local communities. Secondly, we assume that reactors can switch between operating modes without delay and incurred cost. Adding this to the high capacity factors of 90% is an assumption that is advantageous to nuclear. Whether SNRs can operate at such high capacity factors and operate flexibly remains highly uncertain given the lack of precedence in the historical track record of such concepts, see Section 2.2.
@@
Regarding cost assumptions, we show that OCC must be drastically reduced for nuclear power to be competitive with renewable technologies. In our approach, we apply a uniform cost assumption for nuclear power that does not consider cost differences between vendors, financing schemes, or local cost differences that could occur between, for example, projects in Finland and Hungary.
@@
The applied model to compute a cost-efficient share of the different nuclear power plants corresponds to or exceeds the state-of-the-art in energy planning. The high temporal resolution using two-hour blocks, resulting in 4380 steps, exceeds typical resolutions and accurately captures the fluctuations of wind and solar power, which are the main competitors of nuclear power. Inevitable limitations of the model impose a positive bias on nuclear power, so our results constitute an upper bound on the cost-efficient share. Most importantly, we neglect the operational constraints of nuclear power and do not impose restrictions on downtimes or production gradients to keep the underlying optimization problem computationally tractable. Of all considered technologies, these constraints are most relevant for nuclear, but according to previous literature, they are acceptable to ignore, especially if a model includes other short-term flexibility, like batteries or demand-side management [133, 134].
6 Conclusion and policy recommendations
@@
Our analysis shows the various potential non-electrical use cases for nuclear reactors. We show that these applications are limited to individual reactors supplying heat to small localized demand centers. Most heat provision experience has been gathered in Eastern Europe, where a handful of reactors have provided district heat to neighboring towns, and to Japan, where some reactors service desalination plants. Consequentially, the industry’s experience in non-electrical use cases is limited, and it should be critically reflected upon propositions of localized heat provision given the challenge of implementing sufficient emergency planning zones, flexible operations, and the availability of potentially cheaper alternatives.
@@
Regarding potential technologies, we find that as of today, both light and non-light water reactor concepts cannot fulfill the prerequisites defined by the Committee on Laying the Foundation for New and Advanced Nuclear Reactors in the United States et al. [13]. Most notably, our assessment shows that none of the discussed concepts is commercially available. Most concepts are in pre-licensing or licensing stages, and even for those under construction, the remaining challenges of becoming competitive and affordable (yet socially acceptable) remain to be proven.
@@
Despite the lack of empirical data, many projections and assumptions exist regarding the future cost of diverse reactor concepts. The discrepancy between FOAK and NOAK cost estimations is substantial, and the range of cost estimations in a given readiness level is also. Notably, NOAK cost estimations are referenced more often, albeit the exact number "n" of required produced units to reach these costs through economies of multiples remains undefined. Compared to actual ongoing reactor construction projects in OECD countries, the cost assumptions are overly optimistic and could even be criticized as misleading. Furthermore, as cost overruns and construction delays have historically characterized the nuclear power industry, the costs for actual reactor projects are likely to be higher.
Nonetheless, we assess the potential application of light water reactors and SNR concepts in a future decarbonized energy system in Europe in the near future. We run five model calculations assuming variying technolgy availability in each of the three general readiness level scenarios (FOAK, mean NOAK, minimum NOAK for OCC, fuel, and OM cost). For FOAK cost scenarios, the cost-optimizing model builds no nuclear and relies on renewables to provide all necessary energy services. At mean NOAK cost, on average 60% of FOAK cost assumptions, nuclear power plant capacities are added to the grid, primarily for heat provision. At minimum NOAK cost, on average $\frac{1}{4}$ of FOAK cost assumptions, nuclear power plants dominate medium and high process heat generation, primarily due to the construction of HTRs and SFRs, while both high- and low-capacity LWRs are mostly neglected.
Thus, to achieve technological readiness for the widespread implementation of non-light water reactor technologies and potentially reduce expansion costs, development efforts should be focused on a limited number of reactor concepts and focussed on technologies capable of providing high temperature process heat. For policymakers, this implies goal-oriented (financial) and reliable support of projects with the highest degree of success and future applicability. Cost reductions are only possible when a substantial number of reactors of the same design are built without continuous re-designs necessitating additional safety inspections and licensing procedures.
@@
Based on the results described above, one could assume that nuclear power could play a substantial role in a future decarbonized energy system. However, the energy model’s assumptions are skewed positively towards nuclear power. This includes optimistic assumptions of 90% capacity factors, the assumption that reactors can be built at NOAK costs in a few years, full flexibility of reactors and the ability to switch between electricity and heat provision without ramp-up, the neglect of decommissioning and waste management costs, potential safety challenges of non-light water reactor concepts, increased proliferation risks, and the uncertainty regarding actual industry capability to deliver the required capacities.
Consequentially, policymakers must address these unresolved challenges of non-light water reactor concepts. Most notably this refers to the organization of new waste disposal routes—all while most countries have yet to safely dispose of today’s fleets spent fuel—as well as the assessment of proliferation and safety risks. Redesigning licensing procedures to account for the high number of necessary reactor installations and speeding up procedures while accounting for heterogeneous designs and associated risks, might become necessary.
To conclude, given the nuclear industry’s current capabilities, the non-availability of a limited number of tested reactor concepts to attempt the proposed mass production, unproven non-electrical applications on a large scale, unsolved waste issues with novel fuel concepts, potential safety challenges, and obvious social challenges for the large scale implementation of nuclear power, it seems unlikely, from a techno-economic perspective, that non-light water reactors will play a major role in a near-future decarbonized energy system.
Thus, while the here discussed reactor concepts might become available in the coming decades, and could at some point be economically competitive, regulators and policymakers should also focus on implementing available technologies, i.e., renewables and grid capacity expansion coupled with demand-side flexibility and sufficient storage options, to continue ongoing energy system decarbonization efforts necessary today.
@@ Acknowledgments
The research leading to these results has received funding from the Deutsche Forschungsgemeinschaft (DFG) under project number 423336886 and the German Federal Office for the Safety of Nuclear Waste Management (BASE), grant number 4722B50101.
@@ Appendix B Non-Light Water Reactor Technologies
LFRs operate with fast neutrons and thus require no moderator. The coolant is liquid lead or lead-bismuth composite. LFRs are planned to run on MOX fuel with output temperatures of 400 to 620°C. The main challenges for LFR implementation relate to the supply of MOX fuel and required infrastructure, the management of generated polonium-210, corrosion and material erosion risks because fuel rods containing iron, nickel and cobalt would be diluted in the coolant, the sensitivity of a potential reactor to seismic activity due to the high density and mass of the coolant, and finally, inspection and monitoring challenges due to opaque coolant. Historically, LFRs were used for marine propulsion, but in addition to the same potential applications of SFRs, they could provide load-following services due to the theoretically short ramp-up times [63, 64, 60]. Initial development efforts began in 1942 in the U.S. but were soon canceled. The first experimental reactor 27/VT was built in 1959 in the Soviet Union for marine propulsion. In the 1970s, seven Soviet Lira-class submarines were fitted with LFRs, all of which have been out of service since the mid-1990s [43]. Currently, there is no operational LFR. Current concepts are the Belgian MYRRHA (also planned as an accelerator-driven system), the Chinese CLFR-100, CLFR-300 and BLESS, the EU-funded projects ALFRED and ELFR, the LRF-AS-200 under development in Luxemburg, the Russian BREST-OD-300 (under construction) and SVBR-100, the South Korean PEACER, and the Westinghouse-LFR in the U.S. [64, 60, 63].
@@
The term MSRs comprises a multitude of different reactor concepts and designs. The common denominator is the chosen coolant being a type of molten salt, based on lithium or sodium in combination with fluoride or chloride. Furthermore, some concepts run on fast, some on slow neutron spectra. The latter require graphite as moderator. Fuel can be sometimes integrated into the molten salts, and be of various liquid and solid concepts, such as HALEU, uranium-plutonium, uranium-thorium, or plutonium combined with minor actinides. Output temperatures range around 600 to 700°C. In addition to heat and electricity provision, MSRs could function as thermal storage facilities. Current challenges limiting MSR establishment relate to the identification and testing of suitable salt-material combinations and the development of integrated models for safety case assessments. Furthermore, the radioactive salts require the establishment of infrastructure for storage and transport. Additionally, proliferation and safety regulations might have to be adapted [63, 68]. The U.S. Air Force first researched MSRs for aircraft propulsion in the 1950s, followed by the MSRE experimental reactor built at the Oak Ridge National Laboratory (ORNL) in the 1960s. This reactor has been in a long-term enclosure since 1969 [145]. There are no currently operating MSRs, but an array of different concepts are under development. Thermal fluoride-based MSRs with solid fuel are under development in China (TSMR-SF) and the U.S. Here, the ORNL is working on the AHTR and Mark 1 PB-FHR concepts, and Kairos Power is developing the KP-FHR. Thermal fluoride-based MSRs with liquid fuel are under development in Canada (Terrestrial Energy’s IMSR-400), China (TSMR-LF), Denmark (Seaborgy Technology’s MSTR or CMSR concept), Japan (FUJI), and the ORNL’s ThorCon in the US. Two projects, the Russian MOSART and the internationally-funded MSFR (funded by Australia, Canada, the EU, France, Russia, Switzerland, and the U.S.) are concepts for fast fluoride-based MSRs with liquid fuel. Finally, three fast chloride-based MSR concepts with liquid fuel are under development. The projects are the German-designed, recently Canada-based Dual-Fluid reactor, and US-based concepts MSCFR by Elysium, and the SSR-W by Moltex Energy [68, 63].
@@ Appendix C Cost and Technology Data
This section of the appendix provides all the data from the literature analysis. Tables 4, 5, and 6 give the minimum, mean, and maximum OCC and OM cost normalized to USD-2019, as well as the number of retrieved data points, per reactor type. Table 7 shows the factors used to normalize the cost data taken from the literature shown in Tables LABEL:tab:cost_raw_occ, LABEL:tab:cost_raw_om_fix and LABEL:tab:cost_raw_om_var.
@@
S-PRISM SFR 1651 1317 USD/kWe 1996 TCC 4 units ???? NOAK [99]
@@
Appendix D Energy vectors and technologies in the system model
The section gives a more detailed overview of the technology options and final demand categories considered in the model. Since the described features of the model are not novel or specific to this publication, the description is very similar to that of Göke et al. [8]. Overall, the model covers 22 energy carriers and 125 technologies, excluding the nuclear technologies we explicitly added for this paper. To illustrate the resulting complexity, Fig 10 Fig 10. In this graph, the entering edges of technologies refer to input carriers, and the outgoing edges refer to outputs.

### v2

@@ 5 Discussion
The results underline the lack of cost competitiveness of nuclear power compared to other energy generation technologies. They show that the cost-efficient capacity expansion of especially non-light water reactors would require substantial cost reductions that are, as described in Section 3.3.2, not to be expected in the coming years.
@@
However, if the above-mentioned optimistic cost reductions were to be achieved, the results show that non-light water reactor technologies could contribute to energy system decarbonization in Europe in this particular setting (e.g., assuming favorable capacity factors and industry capability, etc.). SFRs and HTRs quickly saturate their heat markets and LWRs are used mainly for electricity production as long as their are cheaper than other (nuclear) technologies. District heat is perceived by the model as useful by product. SMRs compete with other low-carbon technologies for low-temperature heat and are strictly dominated by LWRs for electricity production; they are thus not built in measurable quantities.
@@
But regardless of costs, nuclear reactors feature additional challenges that would need to be overcome for a substantial capacity ramp-up. For example, the heterogeneity of fuel requirements for different SNR concepts corresponds with that of the reactor designs themselves. Many concepts additionally increase proliferation risks because of their requirement of high assay low enriched uranium (HALEU) fuel for which specialized fuel production facilities would have to be funded and constructed, which in turn requires substantial investment and long-term planning and the limitation of industry efforts to a handful of promising reactor designs [152, 153, 17]. It is uncertain whether this type of infrastructure could be up and running in time for these reactor concepts to play a significant role in decarbonizing the European energy system.
@@
Furthermore, the cost data mentioned above do not include back-end costs of nuclear, such as decommissioning and waste management, or social costs, such as risks relating to accidents, proliferation, or uranium mining and tilling [154, 155, 156, 157]. Additional waste challenges arise from implementing non-light water reactors that generate different kinds of waste from, e.g., HALEU fuel, creating additional and currently unstudied waste streams that would have to be dealt with [153]. However, with most nuclear waste responsibility lying with the respective governments [158, 159], substantial funding uncertainties could be removed from projects [160], as suggested to be implemented by Sawicki and Horbaczewska [161]. This would, however, require substantial political power to be achieved, especially in Europe. Additionally, long-term financial assurance and commitment to a limited number of reactor designs are necessary [19].
@@
Additionally, the assumed flexibility of nuclear power plant operations without constraints can be contested. Lynch et al. [88] demonstrate that today’s LWRs can be an effective technology for load following, and Jenkins et al. [22] show in their model for the South-West U.S. that flexible nuclear operations can reduce system costs and increase reactor operator revenues. Loisel et al. [162, p. 281] even claim that "in practice, countries with […] high intermittent renewables […] need [nuclear power plants] to operate". However, despite theoretical technical feasibility, current nuclear reactors’ ability to run flexibly is seldom applied today, with operational experience limited mostly to France [51]. In the U.S., operational experience of flexibly dispatchable nuclear power plants has been limited despite regulations allowing for flexible operations and load following [163]. Reactors designed to operate as flexible backup capacities in energy systems based mainly on renewables face challenges of long standstills and fast ramp-ups. Running at high capacity factors generating heat for other purposes and then flexibly switching to power generation when required will be technologically challenging to implement [164, 64].
@@
Our modelling approach faces some inevitable limitations that could limit the applicability of the results. First, the potential locality of heat demand is not accounted for. This is relevant because, for direct heat provision, sources must be located close to off-takers to minimize losses during transfer. This could be especially relevant for nuclear reactors as these technologies may face substantial opposition from local communities. Secondly, we assume that reactors can switch between operating modes without delay and incurred cost. Adding this to the high capacity factors of 90% is an assumption that is advantageous to nuclear. Whether non-LWR concepts can operate at such high capacity factors and operate flexibly remains highly uncertain given the lack of precedence in the historical track record of such concepts, see Section 2.2.
@@
Regarding cost assumptions, we show that OCC must be drastically reduced for nuclear power to be competitive with renewable technologies. In our approach, we apply a uniform cost assumption for nuclear power that does not consider cost differences between vendors, financing schemes, or local cost differences that could occur between, for example, projects in Finland and Hungary or the U.K.
@@
The applied model to compute a cost-efficient share of the different nuclear power plants corresponds to or exceeds the state-of-the-art in energy planning. Inevitable limitations of the model impose a positive bias on nuclear power, so our results constitute an upper bound on the cost-efficient share. Most importantly, we neglect the operational constraints of nuclear power and do not impose restrictions on downtimes or production gradients to keep the underlying optimization problem computationally tractable. Of all considered technologies, these constraints are most relevant for nuclear, but according to previous literature, they are acceptable to ignore, especially if a model includes other short-term flexibility, like batteries or demand-side management [165, 166].
6 Conclusion
@@
Our literature analysis shows the various potential non-electrical use cases for nuclear reactors and associatated costs for discussed reactor designs. We show that, as of today, these applications are limited to individual reactors supplying heat to small localized demand centers. Most heat provision experience has been gained in Eastern Europe, where a handful of reactors have provided district heat to neighboring towns, and to Japan, where some reactors service desalination plants. Consequentially, the industry’s experience in non-electrical use cases is limited, and propositions of localized heat provision should consider the challenges regarding the implementation of sufficient emergency planning zones, flexible operations, and the availability of potentially cheaper alternatives.
@@
Regarding potential technologies, we find that as of today, both SMR and non-LWR concepts cannot fulfill the prerequisites defined by the Committee on Laying the Foundation for New and Advanced Nuclear Reactors in the United States et al. [14]. We demonstrate that none of the discussed concepts is commercially available. Most concepts are in pre-licensing or licensing stages. However, with ongoing research and development efforts, first concepts might become available towards the turn of the decade. Regardless, the remaining challenges of becoming competitive and affordable (yet socially acceptable) remain to be proven.
@@
Despite the lack of empirical data, many projections and assumptions exist regarding the future cost of diverse reactor concepts. The discrepancy between FOAK and NOAK cost estimations is substantial, and the range of cost estimations in a given readiness level is, too. Notably, NOAK cost estimations are referenced more often, albeit the exact number "n" of required produced units to reach these costs through economies of multiples remains undefined. Compared to ongoing reactor construction projects in OECD countries, the cost assumptions are highly optimistic. Furthermore, as cost overruns and construction delays have historically characterized the nuclear power industry, costs for actual reactor projects are likely to be higher.
Nonetheless, we assess the potential application of LWR and non-LWR concepts in a decarbonized near-future energy system in Europe. In ten scenarios, we continuously reduce OCC values for four different reactor types. We find that nuclear capacities are built primarily for energy types whose provision requires similarly expensive technologies, i.e., medium and high temperature process heat. The required cost ranges seem rather unlikely given the early stages of development for most non-LWR and SMR concepts as well as past experiences from high-capacity LWR construction projects in Europe.
@@
The results described above give an optimistic upper-bound of the potential role of nuclear power in a future energy system because the model’s assumptions are skewed positively towards the considered reactor technologies. This includes optimistic assumptions of 90% capacity factors, the assumption that reactors can be built at NOAK costs within a few years, full operational flexibility of reactors and the ability to switch between electricity and heat provision without ramp-up restrictions, as well as the neglect of decommissioning and waste management costs, potential safety challenges of non-LWR concepts, proliferation risks, and the uncertainty regarding actual industry capability to deliver the required capacities.
Consequentially, policymakers must address these unresolved industrial challenges of non-LWR concepts before a large-scale implementation is possible. Additionally, open questions regarding the organization of new waste disposal routes and potentially increased proliferation and safety risks must be addressed.
@@ Acknowledgments
The authors with to thank the participants of the 2025 IEWT Conference in Vienna, of the 2024 International IAEE Conference in Istanbul and of the 2025 AT-OM Day held in Berlin for helpful comments on earlier versions of this work. We extend our gratitude to Christian von Hirschhausen for his suggestions for the improvement of this work.
Funding
The research leading to these results has received funding from the Deutsche Forschungsgemeinschaft (DFG) under project number 423336886. We thank the University Library of TU Berlin for providing funds for an open access publication.
@@ Appendix B Non-Light Water Reactor Technologies
LFRs operate with fast neutrons and thus require no moderator. The coolant is liquid lead or lead-bismuth composite. LFRs are planned to run on MOX fuel with output temperatures of 400 to 620°C. The main challenges for LFR implementation relate to the supply of MOX fuel and required infrastructure, the management of generated polonium-210, corrosion and material erosion risks because fuel rods containing iron, nickel and cobalt would be diluted in the coolant, the sensitivity of a potential reactor to seismic activity due to the high density and mass of the coolant, and finally, inspection and monitoring challenges due to opaque coolant. Historically, LFRs were used for marine propulsion, but in addition to the same potential applications of SFRs, they could provide load-following services due to the theoretically short ramp-up times [77, 78, 74, 25]. Initial development efforts began in 1942 in the U.S. but were soon canceled. The first experimental reactor 27/VT was built in 1959 in the Soviet Union for marine propulsion. In the 1970s, seven Soviet Lira-class submarines were fitted with LFRs, all of which have been out of service since the mid-1990s [52]. Currently, there is no operational LFR. Current concepts are the Belgian MYRRHA (also planned as an accelerator-driven system), the Chinese CLFR-100, CLFR-300 and BLESS, the EU-funded projects ALFRED and ELFR, the LRF-AS-200 under development in Luxemburg, the Russian BREST-OD-300 (under construction) and SVBR-100, the South Korean PEACER, and the Westinghouse-LFR in the U.S. [78, 74, 77].
@@
The term MSRs comprises a multitude of different reactor concepts and designs. The common denominator is the chosen coolant being a type of molten salt, based on lithium or sodium in combination with fluoride or chloride. Furthermore, some concepts run on fast, some on slow neutron spectra. The latter require graphite as moderator. Fuel can be sometimes integrated into the molten salts, and be of various liquid and solid concepts, such as HALEU, uranium-plutonium, uranium-thorium, or plutonium combined with minor actinides. Output temperatures range around 600 to 700°C. In addition to heat and electricity provision, MSRs could function as thermal storage facilities [25]. Current challenges limiting MSR establishment relate to the identification and testing of suitable salt-material combinations and the development of integrated models for safety case assessments. Furthermore, the radioactive salts require the establishment of infrastructure for storage and transport. Additionally, proliferation and safety regulations might have to be adapted [77, 82]. The U.S. Air Force first researched MSRs for aircraft propulsion in the 1950s, followed by the MSRE experimental reactor built at the Oak Ridge National Laboratory (ORNL) in the 1960s. This reactor has been in a long-term enclosure since 1969 [178]. There are no currently operating MSRs, but an array of different concepts are under development. Thermal fluoride-based MSRs with solid fuel are under development in China (TSMR-SF) and the U.S. Here, the ORNL is working on the AHTR and Mark 1 PB-FHR concepts, and Kairos Power is developing the KP-FHR. Thermal fluoride-based MSRs with liquid fuel are under development in Canada (Terrestrial Energy’s IMSR-400), China (TSMR-LF), Denmark (Seaborgy Technology’s MSTR or CMSR concept), Japan (FUJI), and the ORNL’s ThorCon in the US. Two projects, the Russian MOSART and the internationally-funded MSFR (funded by Australia, Canada, the EU, France, Russia, Switzerland, and the U.S.) are concepts for fast fluoride-based MSRs with liquid fuel. Finally, three fast chloride-based MSR concepts with liquid fuel are under development. The projects are the German-designed, recently Canada-based Dual-Fluid reactor, and US-based concepts MSCFR by Elysium, and the SSR-W by Moltex Energy [82, 77].
@@ Appendix C Cost and Technology Data
This section of the appendix provides all the data from the literature analysis. Tables 3, 4, and 5 give the minimum, mean, and maximum OCC and OM cost normalized to USD-2019, as well as the number of retrieved data points, per reactor type. Table 6 shows the factors used to normalize the cost data taken from the literature shown in Tables LABEL:tab:cost_raw_occ, LABEL:tab:cost_raw_om_fix and LABEL:tab:cost_raw_om_var.
@@
S-PRISM SFR 1651 1317 USD/kWe 1996 TCC 4 units (assumed) NOAK [125]
@@
Generic SMR SMR 150 7097 EUR/kWe 2025 2030 data NL FOAK [137]
Generic SMR SMR 150 6104 EUR/kWe 2025 2050 data NL FOAK [137]
HTGR-200 HTR 2581 USD/kWe 2023 NOAK [138]
CAREM25 LWR-SMR 3600 USD/kWe 2024 NOAK [139]
KLT-40s LWR-SMR 3950 USD/kWe 2024 NOAK [139]
HTR-PM HTR 1500 USD/kWe 2024 NOAK [139]
Oklo Microreactor LWR-SMR 1.5 6700 USD/kWe 2020 Capital cost NOAK [140]
SMR LWR-SMR 3 5000 USD/kWe 2023 Low Capital cost NOAK [140]
SMR LWR-SMR 3 25000 USD/kWe 2023 High Capital cost NOAK [140]
Microreactor LWR-SMR 5 15000 USD/kWe 2019 Nominal microreactor data NOAK [61]
Gen III+ LWR 1000 2600 USD/kWe 2012 Low estimate NOAK [91]
Gen III+ LWR 1000 6600 USD/kWe 2012 High estimate NOAK [91]
45MW SMR LWR-SMR 45 4000 USD/kWe 2012 Low estimate NOAK [91]
45MW SMR LWR-SMR 45 16300 USD/kWe 2012 High estimate NOAK [91]
225 MW SMR LWR-SMR 225 3200 USD/kWe 2012 Low estimate NOAK [91]
225 MW SMR LWR-SMR 225 7100 USD/kWe 2012 High estimate NOAK [91]
SMR LWR-SMR 3000 USD/kWe 2025 Max cost for ERCOT new build NOAK [141]
SMR LWR-SMR 5029 USD/kWe 2014 Nuclear cost metrics, additional fuel cost: 8.126 USD/MWh NOAK [142]
Conventional Nuclear LWR 5452 USD/kWe 2014 Nuclear cost metrics, additional fuel cost: 8.126 USD/MWh NOAK [142]
Molten Salt Reactor MSR 3000 USD/kWe 2014 Nuclear cost metrics, additional fuel cost: 0.381 USD/MWh NOAK [142]
NuScale VOYGR LWR-SMR 77 14575 USD/kWe 2022 Estimated SMR norm. Occ FOAK [143]
NuScale VOYGR LWR-SMR 77 2900 USD/kWe 2022 Estimated SMR norm. Occ NOAK [143]
Holtec SMR-160 LWR-SMR 160 10878 USD/kWe 2022 Estimated SMR norm. Occ FOAK [143]
Holtec SMR-160 LWR-SMR 160 3750 USD/kWe 2022 Estimated SMR norm. Occ NOAK [143]
Westinghouse AP300 LWR-SMR 300 8459 USD/kWe 2022 Estimated SMR norm. Occ FOAK [143]
Westinghouse AP300 LWR-SMR 300 3450 USD/kWe 2022 Estimated SMR norm. Occ NOAK [143]
Rolls-Royce SMR LWR-SMR 470 7069 USD/kWe 2022 Estimated SMR norm. Occ FOAK [143]
Rolls-Royce SMR LWR-SMR 470 5250 USD/kWe 2022 Estimated SMR norm. Occ NOAK [143]
Supercritical LFR LFR 4356 USD/kWe 2023 Capital cost (OCC) NOAK [144]
Lead-cooled fast reactor LFR 4245 USD/kWe 2023 Capital cost (OCC) NOAK [25]
Large reactor LWR 1340 4250 USD/kWe 2020 OCC could cost NOAK [145]
SMR LWR-SMR 335 4505 USD/kWe 2020 OCC could cost NOAK [145]
SMR LWR-SMR 335 5355 USD/kWe 2020 OCC could cost NOAK [145]
Pebble Bed Modular Reactor (PBMR) HTR 165 4922 USD/kWe 2013 Input data for LCOE calculation + fuel cost 9.33 USD/MWh NOAK [146]
Large nuclear LWR 1000 4102 USD/kWe 2013 Input data for LCOE calculation + fuel cost 9.33 USD/MWh NOAK [146]
SMR LWR-SMR 7031 USD/kWe 2024 Average SMR OCC data NOAK [26]
@@
Table 8: OM data from literature analysis for different reactor types and (assumed) readiness levels. Abbreviations: SFR = Sodium-cooled fast reactor; LWR-SMR = Light water small modular reactor; HTR = High-temperature reactor; LWR = High-capacity ligh water reactor; HTGR = High temperature gas-cooled reactor; MSR = Molten salt reactor; SMR = Small modular reactor; NGNP = Next generation nuclear power plant; PWR = Pressurized water reactor
@@
Generic SMR SMR 150 100.4 EUR/kW 2025 2030 data NL FOAK [137]
Generic SMR SMR 150 100.4 EUR/kW 2025 2050 data NL FOAK [137]
Microreactor LWR-SMR 5 350 USD/kW 2019 Nominal microreactor data NOAK [61]
SMR LWR-SMR n.a. 105 USD/kWh 2025 Max cost for ERCOT new build NOAK [141]
SMR LWR-SMR n.a. 94.72 USD/kW 2014 Nuclear cost metrics, additional fuel cost: 8.126 USD/MWh NOAK [142]
Conventional Nuclear LWR n.a. 80 USD/kW 2014 Nuclear cost metrics, additional fuel cost: 8.126 USD/MWh NOAK [142]
Molten Salt Reactor MSR n.a. 90.26 USD/kW 2014 Nuclear cost metrics, additional fuel cost: 0.381 USD/MWh NOAK [142]
@@
Generic SMR SMR 150 17.8 EUR/kWh 2025 2030 data NL FOAK [137]
Generic SMR SMR 150 17.8 EUR/kWh 2025 2050 data NL FOAK [137]
SMR LWR-SMR n.a. 2.17 USD/MWh 2014 Nuclear cost metrics, additional fuel cost: 8.126 USD/MWh NOAK [142]
Conventional Nuclear LWR n.a. 2.17 USD/MWh 2014 Nuclear cost metrics, additional fuel cost: 8.126 USD/MWh NOAK [142]
Molten Salt Reactor MSR n.a. 2.17 USD/MWh 2014 Nuclear cost metrics, additional fuel cost: 0.381 USD/MWh NOAK [142]
Pebble Bed Modular Reactor (PBMR) HTR 165 14.74 USD/MWh 2013 Input data for LCOE calculation + fuel cost 9.33 USD/MWh NOAK [146]
Large nuclear LWR 1000 14.74 USD/MWh 2013 Input data for LCOE calculation + fuel cost 9.33 USD/MWh NOAK [146]
SMR LWR-SMR n.a. 9 USD/MWh 2024 Average SMR OCC data NOAK [26]
@@
Appendix D Visualization of Model Results for Heat Types
We show lowest and highest nuclear cost scenario results from Germany for each heat type. For district heat, we also show the second-to-last cost scenario. For low temperature process heat, we show data from the French grid because the German case does not include any nuclear for this temperature level.
(a) Highest cost scenario (mean FOAK)
(b) Least cost scenario (minimum NOAK)
Figure 9: Exemplary medium-temperature process heat generation in Germany
(a) Highest cost scenario (mean FOAK)
(b) Least cost scenario (minimum NOAK)
Figure 10: Exemplary low-temperature process heat generation in France
(a) Highest cost scenario (mean FOAK)
(b) Second-to-last cost scenario
(c) Highest cost scenario (mean FOAK)
Figure 11: Exemplary district heat generation in Germany
Appendix E Energy Vectors and Technologies in the System Model
The section gives a more detailed overview of the technology options and final demand categories considered in the model. Since the described features of the model are not novel or specific to this publication, the description is very similar to that of goke_flexible_2023. Overall, the model covers 22 energy carriers and 125 technologies, excluding the nuclear technologies we explicitly added for this paper. To illustrate the resulting complexity, Fig 12 Fig 12. In this graph, the entering edges of technologies refer to input carriers, and the outgoing edges refer to outputs.